OpenStudy (anonymous):

The legs of a right triangle measure 9 feet and 12 feet. What is the length of the hypotenuse? A. 21 feet B. 18 feet C. 15 feet D. 225 feet

OpenStudy (anonymous):

you pythagoras theorem \[a^2 + b^2 = c^2\] a = 9 b= 12 c = what you are looking for